# Kinetic Interacting Particle Langevin Monte Carlo

### v2

@@ Appendix A Preliminary results
Let $z=(z_{\theta},z_{1},\ldots,z_{N})$ and $z^{\prime}=(z_{\theta}^{\prime},z_{1}^{\prime},\ldots,z_{N}^{\prime})$. We have
$\langle z-z^{\prime},\nabla\bar{U}_{N}(z)-\nabla\bar{U}_{N}(z^{\prime})\rangle$ $=\left\langle\begin{pmatrix}z_{\theta}-z_{\theta}^{\prime}\\ z_{1}-z_{1}^{\prime}\\ \vdots\\ z_{N}-z_{N}^{\prime}\end{pmatrix},\begin{pmatrix}(1/N)\sum_{i=1}^{N}\left(\nabla_{\theta}U(z_{\theta},\sqrt{N}z_{i})-\nabla_{\theta}U(z_{\theta}^{\prime},\sqrt{N}z_{i}^{\prime})\right)\\ N^{-1/2}\nabla_{x}U(z_{\theta},\sqrt{N}z_{1})-N^{-1/2}\nabla_{x}U(z_{\theta}^{\prime},\sqrt{N}z_{1}^{\prime})\\ \vdots\\ N^{-1/2}\nabla_{x}U(z_{\theta},\sqrt{N}z_{N})-N^{-1/2}\nabla_{x}U(z_{\theta}^{\prime},\sqrt{N}z_{N}^{\prime})\end{pmatrix}\right\rangle$
$=\frac{1}{N}\sum_{i=1}^{N}\left(\langle z_{\theta}-z_{\theta}^{\prime},\nabla_{\theta}U(z_{\theta},\sqrt{N}z_{i})-\nabla_{\theta}U(z_{\theta}^{\prime},\sqrt{N}z_{i}^{\prime})\rangle\right)$
$\quad+\frac{1}{N}\sum_{i=1}^{N}\langle N^{1/2}z_{i}-N^{1/2}z_{i}^{\prime},\nabla_{x}U(z_{\theta},\sqrt{N}z_{i})-\nabla_{x}U(z_{\theta}^{\prime},\sqrt{N}z_{i}^{\prime})\rangle.$
Next, H1 implies that
$\langle z-z^{\prime},\nabla\bar{U}_{N}(z)-\nabla\bar{U}_{N}(z^{\prime})\rangle\geq\frac{1}{N}\sum_{i=1}^{N}\mu\left(\|z_{\theta}-z_{\theta}^{\prime}\|^{2}+N\|z_{i}-z_{i}^{\prime}\|^{2}\right)=\mu\|z-z^{\prime}\|^{2}.$
This completes the proof. $\Box$
@@
$\|\nabla\bar{U}_{N}(z)-\nabla\bar{U}_{N}(z^{\prime})\|^{2}$ $=\left\|\begin{pmatrix}(1/N)\sum_{i=1}^{N}\nabla_{\theta}U(z_{\theta},\sqrt{N}z_{i})-\nabla_{\theta}U(z_{\theta}^{\prime},\sqrt{N}z_{i}^{\prime})\\ N^{-1/2}\nabla_{x}U(z_{\theta},\sqrt{N}z_{1})-N^{-1/2}\nabla_{x}U(z_{\theta}^{\prime},\sqrt{N}z_{1}^{\prime})\\ \vdots\\ N^{-1/2}\nabla_{x}U(z_{\theta},\sqrt{N}z_{N})-N^{-1/2}\nabla_{x}U(z_{\theta}^{\prime},\sqrt{N}z_{N}^{\prime})\end{pmatrix}\right\|^{2}$
$=\left\|\frac{1}{N}\sum_{i=1}^{N}\left(\nabla_{\theta}U(z_{\theta},\sqrt{N}z_{i})-\nabla_{\theta}U(z_{\theta}^{\prime},\sqrt{N}z_{i}^{\prime})\right)\right\|^{2}$
$\quad+\frac{1}{N}\sum_{i=1}^{N}\left\|\nabla_{x}U(z_{\theta},\sqrt{N}z_{i})-\nabla_{x}U(z_{\theta}^{\prime},\sqrt{N}z_{i}^{\prime})\right\|^{2}.$
Using the fact that $(\cdot)^{2}$ is a convex function and Jensen’s inequality for the first term, we obtain
@@
$\quad+\frac{1}{N}\sum_{i=1}^{N}\left\|\nabla_{x}U(z_{\theta},\sqrt{N}z_{i})-\nabla_{x}U(z_{\theta}^{\prime},\sqrt{N}z_{i}^{\prime})\right\|^{2}.$
@@
$\|\nabla\bar{U}_{N}(z)-\nabla\bar{U}_{N}(z^{\prime})\|^{2}$ $\leq\frac{L^{2}}{N}\sum_{i=1}^{N}\left(\|z_{\theta}-z_{\theta}^{\prime}\|^{2}+N\|z_{i}-z_{i}^{\prime}\|^{2}\right)$
$=L^{2}\|z-z^{\prime}\|^{2},$
@@
Appendix B Proofs
B.1 Proof of Proposition 1
Given Lemma A.1, the system (15) has a positive recurrent Markov stationary measure, absolutely continuous with respect to the Lebesgue measure, given by the density [34]
$\bar{\pi}^{N}(z,\bar{v})\propto\exp\left(-N\bar{U}_{N}(z)-\frac{N}{2}\|\bar{v}\|^{2}\right),$
where $\bar{U}_{N}$ is given in (16). Note, this follows from eqs. (1.1) and (1.3) in [34] by putting $U:=N\bar{U}_{N}$ and $u=N^{-1}$. Let $z=(\theta,z_{1},\ldots,z_{N})$. Using (16), we have

### v3

@@ Appendix A Preliminary results
Let
$z=(z_{\theta},z_{1},\dots,z_{N}),\hskip 20.00003ptz^{\prime}=(z_{\theta}^{\prime},z_{1}^{\prime},\dots,z_{N}^{\prime}).$
From
$\bar{U}_{N}(z)=\frac{1}{N}\sum_{i=1}^{N}U(z_{\theta},\sqrt{N}\,z_{i}),$
we have
$\nabla_{z_{\theta}}\bar{U}_{N}(z)=\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}U(z_{\theta},\sqrt{N}z_{i}),\quad\nabla_{z_{i}}\bar{U}_{N}(z)=\frac{1}{\sqrt{N}}\nabla_{x}U(z_{\theta},\sqrt{N}z_{i}).$
Hence
$\langle z-z^{\prime},\nabla\bar{U}_{N}(z)-\nabla\bar{U}_{N}(z^{\prime})\rangle=\frac{1}{N}\sum_{i=1}^{N}\langle z_{\theta}-z_{\theta}^{\prime},\Delta_{\theta,i}\rangle+\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\langle z_{i}-z_{i}^{\prime},\Delta_{x,i}\rangle,$
where
$\Delta_{\theta,i}:=\nabla_{\theta}U(z_{\theta},\sqrt{N}z_{i})-\nabla_{\theta}U(z_{\theta}^{\prime},\sqrt{N}z_{i}^{\prime}),$
$\Delta_{x,i}:=\nabla_{x}U(z_{\theta},\sqrt{N}z_{i})-\nabla_{x}U(z_{\theta}^{\prime},\sqrt{N}z_{i}^{\prime}).$
Now apply Assumption H1 to the two points
$(z_{\theta},\sqrt{N}z_{i}),\hskip 20.00003pt(z_{\theta}^{\prime},\sqrt{N}z_{i}^{\prime}).$
For each $i$,
$\langle z_{\theta}-z_{\theta}^{\prime},\Delta_{\theta,i}\rangle+\sqrt{N}\,\langle z_{i}-z_{i}^{\prime},\Delta_{x,i}\rangle\geq\mu\big(\|z_{\theta}-z_{\theta}^{\prime}\|^{2}+N\|z_{i}-z_{i}^{\prime}\|^{2}\big).$
Divide by $N$ and sum over $i$:
$\langle z-z^{\prime},\nabla\bar{U}_{N}(z)-\nabla\bar{U}_{N}(z^{\prime})\rangle\geq\mu\|z_{\theta}-z_{\theta}^{\prime}\|^{2}+\mu\sum_{i=1}^{N}\|z_{i}-z_{i}^{\prime}\|^{2}=\mu\|z-z^{\prime}\|^{2}.$
So $\bar{U}_{N}$ is $\mu$-strongly convex. $\Box$
@@
$\|\nabla\bar{U}_{N}(z)-\nabla\bar{U}_{N}(z^{\prime})\|^{2}=$ $\left\|\frac{1}{N}\sum_{i=1}^{N}\left(\nabla_{\theta}U(z_{\theta},\sqrt{N}z_{i})-\nabla_{\theta}U(z_{\theta}^{\prime},\sqrt{N}z_{i}^{\prime})\right)\right\|^{2}$
$+\frac{1}{N}\sum_{i=1}^{N}\left\|\nabla_{x}U(z_{\theta},\sqrt{N}z_{i})-\nabla_{x}U(z_{\theta}^{\prime},\sqrt{N}z_{i}^{\prime})\right\|^{2}.$
Using the fact that $\|\cdot\|^{2}$ is a convex function and Jensen’s inequality for the first term, we obtain
@@
$+\frac{1}{N}\sum_{i=1}^{N}\left\|\nabla_{x}U(z_{\theta},\sqrt{N}z_{i})-\nabla_{x}U(z_{\theta}^{\prime},\sqrt{N}z_{i}^{\prime})\right\|^{2}.$
@@
$\|\nabla\bar{U}_{N}(z)-\nabla\bar{U}_{N}(z^{\prime})\|^{2}\leq\frac{L^{2}}{N}\sum_{i=1}^{N}\left(\|z_{\theta}-z_{\theta}^{\prime}\|^{2}+N\|z_{i}-z_{i}^{\prime}\|^{2}\right)=L^{2}\|z-z^{\prime}\|^{2},$
@@
Appendix B Equivalence of algorithms
We prove in Proposition A.1 that the KIPLD can be rewritten as a single underdamped Langevin diffusion evolving in ${\mathbb{R}}^{d_{\theta}+Nd_{x}}$. This brings the natural question about the algorithms we derived, namely, KIPLMC1 and KIPLMC2, which are discretisations of the KIPLD, and how they relate to the standard discretisation schemes the rescaled diffusion in (12). We show below that the standard discretisations of (12) are precisely the KIPLMC1 and KIPLMC2 algorithms we defined in Section 3. This will later allow us to leverage existing results on the convergence of these standard discretisation schemes to obtain convergence results for our algorithms.
B.1 KIMPLC1
We show below the exponential integrator algorithm, presented in Dalalyan and Riou-Durand [19], for our diffusion (KIPLD) matches the KIPLMC1 algorithm we defined in Section 3.
Proposition A.2.
Let the sequence $(\widetilde{Z}_{k},\widetilde{V}_{k}^{z})_{k\geq 0}$ be the exponential-integrator discretisation of the rescaled underdamped diffusion (12) with friction parameter $\widetilde{\gamma}=\sqrt{N}\gamma$ and step-size $\widetilde{\eta}=\eta/\sqrt{N}$. Define the unscaled variables by
$\widetilde{\theta}_{k}$ $=\theta_{k},\qquad\widetilde{X}_{k}^{i}=N^{-1/2}X_{k}^{i},$
$\widetilde{V}_{k}^{\theta}$ $=\sqrt{N}V_{k}^{\theta},\qquad\widetilde{V}_{k}^{x_{i}}=V_{k}^{x_{i}},$
for each $i\in[N]$. Then, after the corresponding rescaling of the Gaussian noise variables, the resulting recursion in $(\theta_{k},X_{k}^{i},V_{k}^{\theta},V_{k}^{x_{i}})_{i\in[N]}$ coincides exactly with the KIPLMC1 scheme defined in Section 3.1.
Proof.
Let us recall that, by Proposition A.1, we can rewrite (KIPLD) as the standard underdamped SDE
${\mathrm{d}}\widetilde{{\mathbf{Z}}}_{t}$ $=\widetilde{{\mathbf{V}}}_{t}^{z}\,{\mathrm{d}}t,$
${\mathrm{d}}\widetilde{{\mathbf{V}}}_{t}^{z}$ $=-\widetilde{\gamma}\,\widetilde{{\mathbf{V}}}_{t}^{z}\,{\mathrm{d}}t-N\nabla_{z}\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{t})\,{\mathrm{d}}t+\sqrt{2\widetilde{\gamma}}\,{\mathrm{d}}\widetilde{{\mathbf{B}}}_{t},$
where $\tilde{\gamma}=\sqrt{N}\gamma$ is the friction parameter and we also implicitly adjust its initial conditioned to be rescaled. We discretise this diffusion using the exponential integrator scheme (termed KLMC1) developed by Dalalyan and Riou-Durand [19]. In order to do this, define a sequence of functions $\tilde{\psi}_{k+1}^{t}=\int_{0}^{t}\tilde{\psi}_{k}^{s}{\mathrm{d}}s$ where $\tilde{\psi}_{0}^{t}=e^{-\tilde{\gamma}t}$. With this definition, the exponential integrator scheme with step-size $\tilde{\eta}$ is defined as
$\widetilde{Z}_{k+1}$ $=\widetilde{Z}_{k}+\tilde{\psi}_{1}^{\tilde{\eta}}\tilde{V}_{k}-\tilde{\psi}_{2}^{\tilde{\eta}}N\nabla_{z}\bar{U}_{N}(\widetilde{Z}_{k})+\sqrt{2\tilde{\gamma}}\zeta_{k+1}^{\prime}$
$\widetilde{V}_{k+1}$ $=\tilde{\psi}_{0}^{\tilde{\eta}}\widetilde{V}_{k}-\tilde{\psi}_{1}^{\tilde{\eta}}N\nabla_{z}\bar{U}_{N}(\tilde{Z}_{k})+\sqrt{2\tilde{\gamma}}\zeta_{k+1},$
where the stacked vector $(\zeta_{k})_{1},\dots,(\zeta_{k})_{d_{z}},(\zeta^{\prime}_{k})_{1},\dots(\zeta^{\prime}_{k})_{d_{z}}]^{\top}$ is a centred Gaussian random variable with covariance $\widetilde{C}\otimes I_{d_{z}}$, where $\widetilde{C}$ is given by
$\widetilde{C}=\int_{0}^{\tilde{\eta}}\begin{bmatrix}\tilde{\psi}_{0}^{t}\\ \tilde{\psi}_{1}^{t}\end{bmatrix}\begin{bmatrix}\tilde{\psi}_{0}^{t}&\tilde{\psi}_{1}^{t}\end{bmatrix}\mathrm{d}t.$
Written explicitly via the definition of $\tilde{\psi}_{i}^{t}$, we have,
$\widetilde{C}=\frac{1}{2\tilde{\gamma}}\begin{pmatrix}(1-e^{-2\tilde{\gamma}\tilde{\eta}})&\frac{(e^{-\tilde{\gamma}\tilde{\eta}}-1)^{2}}{\tilde{\gamma}}\\ \frac{(e^{-\tilde{\gamma}\tilde{\eta}}-1)^{2}}{\tilde{\gamma}}&\frac{2\tilde{\gamma}\tilde{\eta}+4e^{-\tilde{\gamma}\tilde{\eta}}-e^{-2\tilde{\gamma}\tilde{\eta}}-3}{\tilde{\gamma}^{2}}\end{pmatrix}$
Let us now relate the matrix $\widetilde{C}$ to the matrix we have defined in (7) for the KIPLMC1 scheme. Written explicitly with the definitions in Section 3.1, KIPLMC1 instead defines functions $\psi_{0}^{\eta}=e^{-\gamma\eta}$ and $\psi_{i+1}^{\eta}=\int_{0}^{\eta}\psi_{i}^{t}{\mathrm{d}}t$, and the covariance matrix in (7) is given as
$C=\frac{1}{2\gamma}\begin{pmatrix}(1-e^{-2\gamma\eta})&\frac{(e^{-\gamma\eta}-1)^{2}}{\gamma}\\ \frac{(e^{-\gamma\eta}-1)^{2}}{\gamma}&\frac{2\gamma\eta+4e^{-\gamma\eta}-e^{-2\gamma\eta}-3}{\gamma^{2}}\end{pmatrix}$.
Recall that the stacked vector $(\varepsilon_{k})_{1},(\varepsilon^{\prime}_{k})_{1},(\varepsilon_{k})_{2},(\varepsilon^{\prime}_{k})_{2},\dots,((\varepsilon_{k})_{d_{z}},(\varepsilon^{\prime}_{k})_{d_{z}})]^{\top}$ is a centred Gaussian random variable with covariance $C\otimes I_{d_{z}}$ as defined in Section 3.1. Note the relation between $\widetilde{C}$ and $C$ is given by
$\widetilde{C}=\begin{pmatrix}N^{-\frac{1}{2}}C_{11}&N^{-1}C_{12}\\ N^{-1}C_{21}&N^{-\frac{3}{2}}C_{22}\end{pmatrix}$.
Hence, we recover
$\zeta_{k}$ $=(N)^{-\frac{1}{4}}(\varepsilon_{k}^{0},\dots,\varepsilon_{k}^{N})^{\top},$
$\zeta_{k}^{\prime}$ $=(N)^{-\frac{3}{4}}(\varepsilon^{0,\prime}_{k},\dots,\varepsilon_{k}^{N,\prime})^{\top}.$
We first write the full exponential-integrator scheme coordinatewise, without yet substituting the relations between $\tilde{\eta}$ and $\eta$ or between the rescaled and original variables. This gives
$\widetilde{\theta}_{k+1}$ $=\widetilde{\theta}_{k}+\tilde{\psi}_{1}^{\tilde{\eta}}\widetilde{V}^{\theta}_{k}-\tilde{\psi}_{2}^{\tilde{\eta}}N\nabla_{\theta}\bar{U}_{N}(\widetilde{Z}_{k})+\sqrt{2\tilde{\gamma}}(\zeta_{k}^{\prime})^{0},$
$\widetilde{X}_{k+1}^{i}$ $=\widetilde{X}_{k}^{i}+\tilde{\psi}_{1}^{\tilde{\eta}}\widetilde{V}_{k}^{x_{i}}-\tilde{\psi}_{2}^{\tilde{\eta}}N\nabla_{z_{i}}\bar{U}_{N}(\widetilde{Z}_{k})+\sqrt{2\tilde{\gamma}}(\zeta_{k}^{\prime})^{i},$
$\widetilde{V}_{k+1}^{\theta}$ $=\tilde{\psi}_{0}^{\tilde{\eta}}\widetilde{V}_{k}^{\theta}-\tilde{\psi}_{1}^{\tilde{\eta}}N\nabla_{\theta}\bar{U}_{N}(\widetilde{Z}_{k})+\sqrt{2\tilde{\gamma}}(\zeta_{k})^{0},$
$\widetilde{V}_{k+1}^{x_{i}}$ $=\tilde{\psi}_{0}^{\tilde{\eta}}\widetilde{V}_{k}^{x_{i}}-\tilde{\psi}_{1}^{\tilde{\eta}}N\nabla_{z_{i}}\bar{U}_{N}(\widetilde{Z}_{k})+\sqrt{2\tilde{\gamma}}(\zeta_{k})^{i},$
where the noise variables $(\zeta_{k})^{0}$, $(\zeta_{k}^{\prime})^{0}$, $(\zeta_{k})^{i}$ and $(\zeta_{k}^{\prime})^{i}$ are the appropriate coordinates of $\zeta_{k}$ and $\zeta_{k}^{\prime}$ respectively. Now recall from Proposition A.1 that $\tilde{\gamma}=\sqrt{N}\gamma$, and choose $\tilde{\eta}=\eta/\sqrt{N}$. Then
$\tilde{\psi}_{0}^{\tilde{\eta}}=e^{-\gamma\eta}=\psi_{0}^{\eta},\quad\tilde{\psi}_{1}^{\tilde{\eta}}=\frac{1}{\gamma\sqrt{N}}(1-e^{-\gamma\eta})=\frac{\psi_{1}^{\eta}}{\sqrt{N}},\quad\tilde{\psi}_{2}^{\tilde{\eta}}=\frac{1}{\gamma N}(\eta-\frac{1}{\gamma}(1-e^{-\gamma\eta}))=\frac{\psi_{2}^{\eta}}{N}.$
Using also the identities
$N\nabla_{\theta}\bar{U}_{N}(\widetilde{Z}_{k})$ $=\sum_{i=1}^{N}\nabla_{\theta}U(\widetilde{\theta}_{k},\sqrt{N}\widetilde{X}_{k}^{i}),$ (14)
$N\nabla_{z_{i}}\bar{U}_{N}(\widetilde{Z}_{k})$ $=\sqrt{N}\nabla_{x}U(\widetilde{\theta}_{k},\sqrt{N}\widetilde{ZX}_{k}^{i}),$
together with $\widetilde{\theta}_{k}=\theta_{k}$, $\widetilde{X}_{k}^{i}=N^{-1/2}X_{k}^{i}$, $\widetilde{V}_{k}^{\theta}=\sqrt{N}V_{k}^{\theta}$, $\widetilde{V}_{k}^{x_{i}}=V_{k}^{x_{i}}$, $(\zeta_{k})^{0}=N^{-1/4}\varepsilon_{k}^{0}$, $(\zeta_{k}^{\prime})^{0}=N^{-3/4}\varepsilon_{k}^{0,\prime}$, $(\zeta_{k})^{i}=N^{-1/4}\varepsilon_{k}^{i}$, and $(\zeta_{k}^{\prime})^{i}=N^{-3/4}\varepsilon_{k}^{i,\prime}$, we obtain for the $\theta$-coordinates
$\theta_{k+1}$ $=\theta_{k}+\frac{\psi_{1}^{\eta}}{\sqrt{N}}\widetilde{V}^{\theta}_{k}-\frac{\psi_{2}^{\eta}}{N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{k},X_{k}^{i})+\sqrt{\frac{2\gamma}{N}}\varepsilon^{0,\prime}_{k},$
$\widetilde{V}^{\theta}_{k+1}$ $=\psi_{0}^{\eta}\widetilde{V}^{\theta}_{k}-\frac{\psi_{1}^{\eta}}{\sqrt{N}}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{k},X_{k}^{i})+\sqrt{2\gamma}\varepsilon^{0}_{k}.$
Since $\widetilde{V}_{k}^{\theta}=\sqrt{N}V_{k}^{\theta}$, substituting this relation into the first line and dividing the second line by $\sqrt{N}$ yields
$\theta_{k+1}$ $=\theta_{k}+\psi_{1}^{\eta}V^{\theta}_{k}-\frac{\psi_{2}^{\eta}}{N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{k},X_{k}^{i})+\sqrt{\frac{2\gamma}{N}}\varepsilon^{0,\prime}_{k},$
$V^{\theta}_{k+1}$ $=\psi_{0}^{\eta}V^{\theta}_{k}-\frac{\psi_{1}^{\eta}}{N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{k},X_{k}^{i})+\sqrt{\frac{2\gamma}{N}}\varepsilon^{0}_{k}.$
For each particle coordinate $i\in[N]$, the same substitutions give
$\widetilde{X}_{k+1}^{i}$ $=\widetilde{X}_{k}^{i}+\frac{\psi_{1}^{\eta}}{\sqrt{N}}V^{x_{i}}_{k}-\frac{\psi_{2}^{\eta}}{\sqrt{N}}\nabla_{x}U(\theta_{k},X_{k}^{i})+\sqrt{\frac{2\gamma}{N}}\varepsilon^{i,\prime}_{k},$
$V^{x_{i}}_{k+1}$ $=\psi_{0}^{\eta}V^{x_{i}}_{k}-\psi_{1}^{\eta}\nabla_{x}U(\theta_{k},X_{k}^{i})+\sqrt{2\gamma}\varepsilon^{i}_{k}.$
Finally, multiplying the first equation by $\sqrt{N}$, i.e. using $X_{k}^{i}=\sqrt{N}\widetilde{X}_{k}^{i}$, yields
$X^{i}_{k+1}$ $=X^{i}_{k}+\psi_{1}^{\eta}V^{x_{i}}_{k}-\psi_{2}^{\eta}\nabla_{x}U(\theta_{k},X_{k}^{i})+\sqrt{2\gamma}\varepsilon^{i,\prime}_{k}$
$V^{x_{i}}_{k+1}$ $=\psi_{0}^{\eta}V^{x_{i}}_{k}-\psi_{1}^{\eta}\nabla_{x}U(\theta_{k},X_{k}^{i})+\sqrt{2\gamma}\varepsilon^{i}_{k}.$
Together with the $\theta$-update above, this is precisely the KIPLMC1 scheme. $\Box$
B.2 KIPLMC2
Proposition A.3.
Let $(\widetilde{Z}_{k},\widetilde{V}_{k}^{z})_{k\geq 0}$ be the OBABO discretisation of the rescaled underdamped diffusion (12) with friction parameter $\widetilde{\gamma}=\sqrt{N}\gamma$ and step-size $\widetilde{\eta}=\eta/\sqrt{N}$. Define the unscaled variables by
$\widetilde{\theta}_{k}$ $=\theta_{k},\qquad\widetilde{X}_{k}^{i}=N^{-1/2}X_{k}^{i},$
$\widetilde{V}_{k}^{\theta}$ $=\sqrt{N}V_{k}^{\theta},\qquad\widetilde{V}_{k}^{x_{i}}=V_{k}^{x_{i}},$
for each $i\in[N]$. Then the resulting recursion in $(\theta_{k},X_{k}^{i},V_{k}^{\theta},V_{k}^{x_{i}})_{i\in[N]}$ coincides exactly with the KIPLMC2 scheme defined in Algorithm 2.
Proof.
In this case we again apply Proposition A.1 to apply the OBABO results from Monmarché [41] for the standard underdamped diffusion, to our own. Let us recall the standard form, given in Proposition A.1 with $\tilde{\gamma}=\sqrt{N}\gamma$,
${\mathrm{d}}\widetilde{{\mathbf{Z}}}_{t}$ $=\widetilde{{\mathbf{V}}}_{t}^{z}\,{\mathrm{d}}t,$
${\mathrm{d}}\widetilde{{\mathbf{V}}}_{t}^{z}$ $=-\widetilde{\gamma}\,\widetilde{{\mathbf{V}}}_{t}^{z}\,{\mathrm{d}}t-N\nabla_{z}\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{t})\,{\mathrm{d}}t+\sqrt{2\widetilde{\gamma}}\,{\mathrm{d}}\widetilde{{\mathbf{B}}}_{t},$
This diffusion is discretised with the OBABO scheme with step-size $\tilde{\eta}$, to obtain,